"find the fraction equivalent to the fraction a/b+b/a-c/d=" ?
Please show me the steps on solving this problem.

Answer by richwmiller(17219)

multiply a/b by ad/ad
multiply b/a by bd/bd
multiply -c/d by ab/ab
a/b+b/a-c/d
a/b*ad/ad=a^2d/abd
b/a by bd/bd=b^2d/abd
-c/d by ab/ab -abc/abd
a^2d/abd+b^2d/abd-abc/abd
now combine them over the same denominator
(a^2d+b^2d-abc)/abd